# Zonal vs. Nodal Pricing: An Analysis of Different Pricing Rules in the German Day-Ahead Market

### v1

@@ 4.1.1 Welfare losses with Euphemia - Impact of Non-Uniform Pricing
Euphemia maximizes welfare subject to linear and anonymous prices, which is different from unconstrained welfare maximization. In our analysis, the average daily welfare loss associated with Euphemia compared to the national configuration (without any network constraints) is 0.13% – or EUR 40,587.32 in absolute terms. This is comparable with previous estimates of 0.05% as a relative loss (Meeus et al.,, 2009), and with an absolute loss of EUR 130,000 for the entire SDAC region reported by All NEMO Committee, 2023a . Disregarding market coupling, the cuts introduced by the Euphemia algorithm to get to linear prices do not deteriorate the outcome by a lot. After redispatch, Euphemia suffers an average daily loss of 0.18% in total costs (or EUR 60,942.65) compared to the national allocation rule and a loss of 1.49% (or EUR 525,587.01) compared to the nodal allocation rule. We argue that the national, zonal, or Euphemia models differ only marginally in terms of allocation, and only a transition to a nodal allocation rule would substantially affect the average generation and redispatch costs.
@@ 4.2 Price Levels
Average CH prices are slightly lower compared to the other pricing rules under consideration, with lower standard deviation and irrespective of the allocation rule. In contrast, median prices are at comparable levels across all pricing rules. Since CH prices are unaffected by pre-determined generator commitments, they tend to be lower and more stable. In particular, prices can be set by a cheaper – yet uncommitted – generator (Schiro et al.,, 2016). In contrast, average Euphemia prices tend to be slightly higher than zonal prices, while median prices are almost equal. This finding implies that the price-setting generator is usually identical for Euphemia and zonal prices, yet sometimes the Euphemia algorithm introduces a cut that disallows a generator that would otherwise be dispatched. As a result, a higher-price generator may set the price.
Price distributions are generally similar across the different zonal configurations for a specific pricing rule. Due to the simplified representation of the transmission network, cross-zonal flow constraints are rarely tight and thus zonal prices are identical to national (single-zone) prices in many hours. This observation suggests that splits into only a few zones, on average, have little impact on prices and thus provide few locational incentives. Table 6 illustrates that, on average, no major price discrepancies between zones can be expected. The zones are depicted in Figure 1.
@@
Zone 1 42.68 (19.85) 41.91 (19.47) 42.19 (19.64) 41.41 (19.05) 42.80 (20.31)
Zone 2 43.17 (19.93) 41.88 (19.16) 42.75 (19.81) 44.45 (20.66)
Zone 3 42.46 (19.95) 42.97 (19.60)
Zone 4 42.99 (19.56)
@@
Figure 1: Zone Configurations
A more in-depth examination of zonal pricing reveals the impact of cross-zonal flows in equalizing prices across different zones. Specifically, while southern zones face a relative shortage of generation capacity compared to demand, the capacity of cross-zonal transmission lines is typically high enough to transmit electricity from northern zones. Our sensitivity analysis revealed that if cross-zonal flows were limited or prohibited, zonal prices can display differences of more than 30 EUR/MWh between northern and southern regions on some days. The small differences in zonal prices observed in Table 6 are due to the cross-zonal transmission capacity (calculated as the sum of thermal limits of all cross-zonal lines) that equalizes prices across zones.
Under nodal pricing (ca. 40,000 prices per day for 1670 nodes – compared to, e.g., 24 national prices per day) more price outliers were observed than under zonal pricing. For instance, 0.08% of IP prices had to be capped at EUR 100 per MWh under the zonal configurations, compared to 2.26% under nodal pricing. Without capping, the maximum observed IP and Join price would be as high as EUR 70,000 per MWh for an hour of extreme scarcity. Note that SDAC also imposes a price cap of EUR 5,000 per MWh for such periods. CH prices, in contrast, are less prone to outliers, peaking at EUR 416 per MWh. Thus tight flow constraints can increase price levels in the short run. The zonal allocation rules oversimplify the transmission network to an extent that nodal prices can hardly be lower, even without congestion. In contrast, when congestion is present in parts of the grid, nodal prices will be higher, resulting in higher average and median prices. Upon further analysis, it was found that nodal prices have high sensitivity to transmission line susceptances, a phenomenon previously described in Bichler & Knörr, (2023).
Figure 2 illustrates nodal IP, CH, and Join prices, averaged over all hours and sorted in ascending order. The dashed horizontal and vertical lines mark the 5th and 95th percentiles for prices and nodes, respectively. Additionally, the figure shows the average national prices and the national average including a uniform EUR/MWh price adder for redispatch costs. This price adder is calculated as the total redispatch costs under min-volume redispatch (ca. EUR 459.0 million) divided by the total demand (ca. 83.3 TWh), and equals 5.51 EUR/MWh. Notably, with redispatch taken into account, higher average prices are expected at some of the nodes. For the majority of nodes, however, the change in average prices is minor or even negative under nodal pricing.
@@
Nodal prices typically vary across locations for each hour, and thereby set locational incentives for flexible units or storage. As seen in Figure 4, standard deviations of nodal prices resemble each other across nodes and pricing rules. Figure 5 illustrates the distribution of hourly standard deviations of prices across nodes, with mean and median standard deviations as solid and dashed lines, respectively. The nodal price variances can be substantial. For example, in the extreme first hour of 2009/11/24, IP prices vary across the country between EUR/MWh 3.45 and EUR/MWh 100. This sets incentives for demand response and short-term adjustments in particular in energy-intensive industrial production. Persistent high nodal prices signal transmission bottlenecks and direct investments in grid expansion.
@@
A primary goal in the design of new bidding zones is the reduction of price variance within a zone (ACER, 2022b, ), which signals congestion. Tables 5 and 6 indicate that the standard deviation of zonal prices does not decrease with the number of zones. However, varying prices can have two main reasons: congestion and temporal changes in supply and demand. Figure 5 considers hourly standard deviations of prices across all nodes, meaning it illustrates the price variance based on grid congestions in each hour, which we refer to as congestion-based standard deviation. In contrast, Figure 6 includes histograms of the standard deviation of nodal prices across all hours, meaning it illustrates the price variance based on temporal effects of supply and demand at each node, which we refer to as time-based standard deviation. Together, Figures 5 and 6 represent a composition of the standard deviations reported for nodal prices in Table 5. Note that prices at a single node can vary substantially over time due to changes in supply and demand. This time-based standard deviation exceeds the congestion-based standard deviation. For example, IP prices exhibit a mean congestion-based standard deviation of EUR/MWh 12.09, compared to a mean time-based standard deviation of EUR/MWh 16.95. These numbers provide evidence that temporal fluctuations of supply and demand have greater impact on price variance compared to congestion.
@@
Zone 1 12.09 12.29 12.31 11.51 12.52
Zone 2 10.45 10.53 12.47 10.97
Zone 3 6.42 6.44
Zone 4 12.85
@@
Zone 1 16.95 16.92 17.04 16.46 17.47
Zone 2 16.99 16.83 17.40 17.22
Zone 3 16.86 16.70
Zone 4 16.11
@@
4.3 GLOCs, LLOCs, MWPs
The average daily GLOCs, LLOCs, and MWPs are summarized in Tables 911, respectively.
@@ 4.3 GLOCs, LLOCs, MWPs
IP 2,743,625.97 2,715,067.90 2,581,180.46 2,665,434.36 3,073,576.90 6,667,208.89
CH 141,338.05 141,785.33 142,372.81 153,430.35 145,411.76 361,031.42
Join 2,784,666.32 2,773,482.10 2,680,900.33 2,719,728.69 3,115,596.04 4,665,204.10
Euphemia 2,772,814.13
Table 9: Average Daily GLOCs
in EUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
IP 0.00 0.00 0.00 0.00 0.00 0.00
CH 33,683.28 32,956.96 33,554.87 40,266.34 35,536.04 179,887.45
Join 989.05 812.56 1,547.75 1,795.06 1,057.30 41,476.85
Euphemia 0.00
Table 10: Average Daily LLOCs
in EUR National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones Nodal
IP 19,812.81 25,662.91 21,626.78 28,348.30 21,558.54 338,926.64
CH 37,345.65 37,658.44 37,975.21 45,054.94 39,156.73 102,818.68
Join 766.2 1,308.67 1,446.57 1,094.99 859.76 23,840.19
@@
By definition, CH prices minimize GLOCs, and on average, they are substantially lower compared to both IP and Join prices, irrespective of the chosen allocation rule. Similar orders of magnitude for GLOCs are observed across all zonal allocation rules, including Euphemia. In contrast, nodal pricing results in the highest GLOCs, aligning with the higher average prices presented in Table 3. Higher prices imply greater forgone profits for generators and increase GLOCs. Generators in nodal markets face stronger incentives to deviate from the optimal outcome, necessitating penalties – a common practice in many U.S. markets (Bichler et al.,, 2022).
As a natural subset of GLOCs, LLOCs follow a similar pattern, with the highest values occurring under nodal pricing, once again correlated with higher average prices. IP pricing ensures zero LLOCs, as discussed in Section 2.3.2, ensuring reliable congestion signals where price differences between nodes only arise during network congestion. In this case, the price difference precisely reflects the marginal value of transmission capacity. In contrast, CH pricing consistently results in the highest LLOCs, supporting observations that congestion signals may be flawed (Schiro et al.,, 2016). This is especially relevant under a nodal allocation rule considering all transmission lines. The Join pricing rule yields substantially fewer LLOCs than CH pricing, indicating a superior quality of congestion signals.
In terms of average MWPs to compensate generator losses, only the Euphemia algorithm generates zero MWPs, albeit at the expense of welfare losses. An interesting observation is that the welfare losses of Euphemia – as discussed in Section 4.1 – exceed, on average, the MWPs required under a national welfare-maximizing allocation, irrespective of the pricing rule. This implies that the additional generated welfare could potentially cover the MWPs under a national allocation, resulting in a net welfare gain. Under nodal pricing, MWPs tend to increase compared to zonal pricing, particularly for IP pricing. This can be attributed to the fact that under a national allocation rule, there exists a single hourly IP price set at the variable cost of the marginal generator. With positive fixed costs and constant variable costs, such a generator will inevitably incur a loss and necessitate MWPs under IP pricing. In contrast, nodal pricing involves several hourly IP prices and price-setting generators across the network, leading to higher overall MWPs. Conversely, the Join pricing rule generates lower MWPs than IP and CH pricing, constituting a negligible share of the total costs outlined in Table 2.
In summary, GLOCs, LLOCs, and MWPs follow similar patterns across different zonal configurations, but increase with nodal pricing. Concerning pricing rules, IP, CH, and Euphemia each minimize a specific class of lost opportunity costs. However, the Join pricing rule stands out by striking a remarkable balance between low MWPs and LLOCs. At the expense of slightly increased GLOCs, Join prices facilitate marginal side-payments and maintain a high quality of congestion signals.
@@ 4.4 Analysis of Representative Days
To corroborate our findings, we examine three selected days in greater detail. Specifically, we analyze November 23 for its low price levels, March 11 for medium price levels, and February 18 for high price levels.
@@
2009/11/23 26,672.26 26,672.26 26,672.26 26,672.26 26,672.26 29,593.76
2009/03/11 30,494.25 30,507.06 30,502.60 30,507.54 30,496.97 33,068.16
2009/02/18 45,481.38 45,483.18 45,483.04 45,666.68 45,486.15 48,718.75
Table 12: Generation Costs: Nov 25, Mar 11, Feb 18
@@
2009/11/23 3,457.50 3,457.50 3,457.50 3,457.50 3,457.50 0.00
2009/03/11 2,969.30 2,989.12 2,955.14 2,965.60 2,990.37 0.00
2009/02/18 3,649.39 3,664.49 3,639.52 3,561.45 3,596.63 0.00
Table 13: Min-Cost Redispatch Costs: Nov 25, Mar 11, Feb 18
@@
On 2009/11/23 and 2009/03/11, the allocation obtained from the Euphemia algorithm is identical to the national configuration, and no welfare loss occurs. On 2009/02/18, costs under the Euphemia algorithm are slightly higher by 0.36%. This validates our assertion that differences between zonal models – including Euphemia – are minor and nuanced, while nodal pricing constitutes a more substantial change in generation and redispatch costs.
@@
2009/11/23 30,129.75 30,129.75 30,129.75 30,129.75 30,129.75 29,593.76
2009/03/11 33,463.55 33,496.18 33,457.73 33,473.14 33,487.34 33,068.16
2009/02/18 49,130.77 49,147.67 49,122.56 49,228.13 49,082.78 48,718.75
Table 14: Total Costs: Nov 25, Mar 11, Feb 18
Figure 7 illustrates the locational IP, CH, and Join prices for the three selected days, averaged over 24 hours. As previously discussed, the prices on these days exhibit a positive correlation with generation costs, with the exception of Northern Central and Southern Germany with consistent price outliers due to local supply structures. Notably, even though at lower price levels, a more distinct north-south price gradient can be observed on 2009/11/23 compared to 2009/03/11. Specifically, on 2009/11/23, demand is low nationwide, yet inexpensive electricity from the north cannot be transmitted to the south, leading to uneven price levels. In contrast, on 2009/03/11, with higher demand but a balanced grid, prices are uniform across the country. This phenomenon holds for all three pricing rules. The simplified zonal models fail to detect any transmission bottlenecks on 2009/11/23, i.e., cross-zonal constraints are not tight and there are no price variations between zones. Consequently, if the zone configurations are suboptimal, zonal prices do not signal scarcity appropriately. B includes an analysis of hourly prices observed on the three days.
@@
(d) 2009/03/11 - Avg IP Prices
(e) 2009/03/11 - Avg CH Prices
(f) 2009/03/11 - Avg Join Prices
@@
Table 8 summarizes the GLOCs, LLOCs, and MWPs for each selected day. On the low-demand day of 2009/11/23, a Walrasian equilibrium as in Definiton 1 is attainable for the zonal models, resulting in zero GLOCs, LLOC, and MWPs. Crucially, all three pricing rules successfully obtain these equilibrium prices. The Walrasian equilibrium, however, can only persist before redispatch is considered and ceases to exist when nodal transmission constraints are taken into account.
GLOCs, LLOCs, and MWPs do not necessarily increase with average price levels. For nodal pricing, they remain relatively consistent across all days. There is a substantial increase for zonal pricing on the medium-price day of 2009/03/11, compared to the high-price day of 2009/02/18. This indicates that incentives to deviate hinge more on the quality of the price signal than the actual price level.
The simplification inherent in zonal models, which leads to redispatch, does not necessarily imply better economic properties of prices. For example, on 2009/03/11 and 2009/02/18, zonal IP and Join prices result in higher GLOCs than nodal pricing. Specifically, a few zonal prices are insufficient to incentivize many market participants to adhere to the allocation.
As discussed above, CH prices always minimize GLOCs, and IP prices always yield zero LLOCs. Although GLOCs are reduced by CH pricing compared to both IP and Join pricing, their LLOCs, and often MWPs, experience a slight increase. In practical terms, this implies that congestion signals are more distorted and higher side-payments are required. In contrast, the Join pricing rule achieves a favorable trade-off between LLOCs and MWPs, requiring only slightly higher penalties (corresponding to GLOCs) than IP pricing.
in kEUR IP CH Join
National GLOC 0.00 0.00 0.00
LLOC 0.00 0.00 0.00
MWP 0.00 0.00 0.00
2 Zones (k) GLOC 0.00 0.00 0.00
LLOC 0.00 0.00 0.00
MWP 0.00 0.00 0.00
2 Zones (s) GLOC 0.00 0.00 0.00
LLOC 0.00 0.00 0.00
MWP 0.00 0.00 0.00
3 Zones GLOC 0.00 0.00 0.00
LLOC 0.00 0.00 0.00
MWP 0.00 0.00 0.00
4 Zones GLOC 0.00 0.00 0.00
LLOC 0.00 0.00 0.00
MWP 0.00 0.00 0.00
Nodal GLOC 1,277.13 272.47 1,492.66
LLOC 0.00 147.44 26.79
MWP 167.32 94.76 34.70
(a) GLOC, LLOC, MWP for 2009/11/23
in kEUR IP CH Join
National GLOC 4,323.93 43.73 4,323.93
LLOC 0.00 10.68 0.00
MWP 0.00 9.04 0.00
2 Zones (k) GLOC 8,009.06 56.54 8,009.06
LLOC 0.00 23.49 0.00
MWP 0.00 23.03 0.00
2 Zones (s) GLOC 5,419.69 52.08 5,419.69
LLOC 0.00 20.17 0.00
MWP 0.00 17.05 0.00
3 Zones GLOC 7,097.32 57.03 7,097.32
LLOC 0.00 23.98 0.00
MWP 0.00 23.53 0.00
4 Zones GLOC 5,152.01 46.45 5,152.01
LLOC 0.00 13.35 0.00
MWP 0.00 11.10 0.00
Nodal GLOC 1,146.96 178.49 1,159.88
LLOC 0.00 96.05 37.93
MWP 94.90 29.73 22.33
(b) GLOC, LLOC, MWP for 2009/03/11
in kEUR IP CH Join
National GLOC 1,157.36 290.42 1,177.70
LLOC 0.00 30.30 9.38
MWP 59.64 22.53 9.29
2 Zones (k) GLOC 1,798.28 288.22 1,772.59
LLOC 0.00 28.16 9.58
MWP 132.61 18.85 9.50
2 Zones (s) GLOC 1,430.45 285.19 1,455.50
LLOC 0.00 25.13 9.36
MWP 58.69 13.73 9.27
3 Zones GLOC 2,387.52 471.72 2,493.28
LLOC 0.00 170.36 9.20
MWP 27.32 164.81 9.02
4 Zones GLOC 1,778.31 291.19 1,860.59
LLOC 0.00 30.96 9.41
MWP 61.81 20.96 9.33
Nodal GLOC 1,024.32 274.54 1,495.84
LLOC 0.00 86.48 14.44
MWP 72.94 34.68 9.78
(c) GLOC, LLOC, MWP for 2009/02/18
Figure 8: Daily GLOCs, LLOCs, MWPs

### v4

@@ 4.1.1 Welfare losses with Euphemia - Impact of Non-Uniform Pricing
Euphemia maximizes welfare subject to linear and anonymous prices, which is different from unconstrained welfare maximization. In our analysis, the average daily welfare loss associated with Euphemia compared to the national configuration (without any network constraints) is 0.34% – or EUR 109,801.24 in absolute terms. This is comparable with previous estimate of 0.05% as a relative loss (Meeus et al.,, 2009). Disregarding market coupling, the cuts introduced by the Euphemia algorithm to get to linear prices do not deteriorate the outcome by a lot. After redispatch, Euphemia suffers an average daily loss of 9.2% (or kEUR 3,683.24) in total costs compared to the nodal allocation rule. We argue that the national, zonal, or Euphemia models differ only marginally in terms of allocation, and only a transition to a nodal allocation rule would substantially affect the average generation and redispatch costs.
@@ 4.2 Price Levels
Average CH prices are slightly lower compared to the other pricing rules under consideration, with lower standard deviation and irrespective of the allocation rule. In contrast, median prices are at comparable levels across all pricing rules. Since CH prices are unaffected by pre-determined generator commitments, they tend to be lower and more stable. In particular, prices can be set by a cheaper – yet uncommitted – generator (Schiro et al.,, 2016). In contrast, average Euphemia prices tend to be slightly higher than zonal prices, while median prices are in the same ranges. This finding implies that the price-setting generator is usually identical for Euphemia and zonal prices, yet sometimes the Euphemia algorithm introduces a cut that disallows a generator that would otherwise be dispatched. As a result, a higher-price generator may set the price.
Price distributions are generally similar across the different zonal configurations for a specific pricing rule. Cross-zonal flow constraints are rarely tight and thus zonal prices are identical to national (single-zone) prices in many hours. This observation suggests that splits into only a few zones, on average, have little impact on prices and thus provide few locational incentives. Table 6 illustrates that, on average, no major price discrepancies between zones can be expected. The zones are depicted in Figure 1.
@@
Zone 1 45.92 (19.76) 46.88 (19.69) 48.36 (19.87) 45.42 (19.68) 42.43 (19.22)
Zone 2 44.02 (19.30) 44.76 (19.55) 44.28 (19.37) 45.58 (19.96)
Zone 3 45.02 (19.28) 44.08 (19.05)
Zone 4 45.38 (19.86)
@@
Figure 1: Zonal Configurations
A more in-depth examination of zonal pricing reveals the impact of cross-zonal flows in equalizing prices across different zones. Specifically, while southern zones face a relative shortage of generation capacity compared to demand, the capacity of cross-zonal transmission lines is typically high enough to transmit electricity from northern zones. Our sensitivity analysis revealed that if cross-zonal flows were limited or prohibited, zonal prices can display differences of more than 30 EUR/MWh between northern and southern regions on some days. The small differences in zonal prices observed in Table 6 are due to the cross-zonal transmission capacity. As described in Section 2.2.1, to compute a zonal solution, we model the interconnectors individually but we do not consider critical network elements and loop flows, which might decrease the available capacity between zones in practice.
Under nodal pricing (ca. 40,000 prices per day for 1670 nodes – compared to, e.g., 24 national prices per day) more price outliers were observed than under zonal pricing. For instance, 0.13% of IP prices had to be capped 100 EUR/MWh under the zonal configurations, compared to 0.25% under nodal pricing. Without capping, the maximum observed IP and Join price would be as high as 62,000 EUR/MWh for an hour of extreme scarcity. Note that SDAC also imposes a price cap of 5,000 EUR/MWh for such periods. CH prices, in contrast, are less prone to outliers, peaking at 423 EUR/MWh. Thus tight flow constraints can increase price levels in the short run. The zonal allocation rules oversimplify the transmission network to an extent that nodal prices can hardly be lower, even without congestion. In contrast, when congestion is present in parts of the grid, nodal prices will be higher, resulting in higher average and median prices. Upon further analysis, it was found that nodal prices have high sensitivity to transmission line susceptances, a phenomenon previously described in Bichler & Knörr, (2023).
Figure 2 illustrates nodal IP, CH, and Join prices, averaged over all hours and sorted in ascending order. The dashed horizontal and vertical lines mark the 5th and 95th percentiles for prices and nodes, respectively. Additionally, the figure shows the average national prices and the national average including a uniform EUR/MWh price adder for redispatch costs. This price adder is calculated as the total redispatch costs under min-volume redispatch (ca. EUR 310.74 million) divided by the total demand (ca. 83.3 TWh), and equals 3.73 EUR/MWh. Notably, with redispatch taken into account, higher average prices are expected at some of the nodes. For the majority of nodes, however, the change in average prices is minor or even negative under nodal pricing.
@@
Nodal prices typically vary across locations for each hour, and thereby set locational incentives for flexible units or storage. As seen in Figure 4, standard deviations of nodal prices resemble each other across nodes and pricing rules. Figure 5 illustrates the distribution of hourly standard deviations of prices across nodes, with mean and median standard deviations as solid and dashed lines, respectively. The nodal price variances can be substantial. For example, in the extreme first hour of 2009/11/23, IP prices vary across the country between 0 EUR/MWh and 100 EUR/MWh. This sets incentives for demand response and short-term adjustments in particular in energy-intensive industrial production. Persistent high nodal prices signal transmission bottlenecks and direct investments in grid expansion.
@@
A primary goal in the design of new bidding zones is the reduction of price variance within a zone (ACER, 2022b, ), which signals congestion. Tables 5 and 6 indicate that the standard deviation of zonal prices does not decrease with the number of zones. However, varying prices can have two main reasons: congestion and temporal changes in supply and demand. Figure 5 considers hourly standard deviations of prices across all nodes, meaning it illustrates the price variance based on grid congestions in each hour, which we refer to as congestion-based standard deviation. In contrast, Figure 6 includes histograms of the standard deviation of nodal prices across all hours, meaning it illustrates the price variance based on temporal effects of supply and demand at each node, which we refer to as time-based standard deviation. Together, Figures 5 and 6 represent a composition of the standard deviations reported for nodal prices in Table 5. Note that prices at a single node can vary substantially over time due to changes in supply and demand. This time-based standard deviation exceeds the congestion-based standard deviation. For example, IP prices exhibit a mean congestion-based standard deviation of EUR/MWh 12.68, compared to a mean time-based standard deviation of EUR/MWh 17.39. These numbers provide evidence that temporal fluctuations of supply and demand have greater impact on price variance compared to congestion.
@@
Zone 1 12.192 12.77 12.71 11.99 12.25
Zone 2 10.25 10.26 12.32 11.16
Zone 3 6.27 6.07
Zone 4 13.21
@@
Zone 1 17.39 17.09 17.39 16.58 18.28
Zone 2 17.65 17.40 18.21 17.56
Zone 3 17.00 17.20
Zone 4 16.12
@@
4.3 MWPs
The average daily MWPs are summarized in Table 9.
@@ 4.3 MWPs
IP 32,400.08 45,448.71 30,881.3 21,122.54 42,156.28 202,106.42
CH 8,691.04 38,029.2 78,808.99 19,794.25 161,948.7 114,669.73
Join 1,153.58 7,375.58 1,780.6 797.12 1,364.48 22,105.27
@@
We observe that only the Euphemia algorithm generates zero MWPs, albeit at the expense of welfare losses. An interesting observation is that the welfare losses of Euphemia – as discussed in Section 4.1 – exceed, on average, the MWPs required under a national welfare-maximizing allocation, irrespective of the pricing rule. This implies that the additional generated welfare could potentially cover the MWPs under a national allocation, resulting in a net welfare gain. Under nodal pricing, MWPs tend to increase compared to zonal pricing, particularly for IP pricing. This can be attributed to the fact that under a national allocation rule, there exists a single hourly IP price set at the variable cost of the marginal generator. With positive fixed costs and constant variable costs, such a generator will inevitably incur a loss and necessitate MWPs under IP pricing. In contrast, nodal pricing involves several hourly IP prices and price-setting generators across the network, leading to higher overall MWPs. Conversely, the Join pricing rule generates lower MWPs than IP and CH pricing, constituting a negligible share of the total costs outlined in Table 2. MWPs follow similar patterns across different zonal configurations, but increase with nodal pricing.
@@ 4.4 Analysis of Representative Days
To corroborate our findings, we examine three selected days in greater detail. Specifically, we analyze November 23 for its low price levels, March 10 for medium price levels, and February 18 for high price levels.
@@
2009/11/23 26,673.74 26,673.74 27,088.39 26,673.74 26,818.18 29,596.10
2009/03/10 31,401.65 31,494.31 31,494.38 31,661.73 31,647.02 35,200.69
2009/02/18 46,287.23 46,471.33 46,287,57 46,287.34 46,432,68 49,303.25
Table 10: Generation Costs: Nov 23, Mar 10, Feb 18
@@
2009/11/23 5,060.40 5,060.40 4,815.01 5,057.49 4,809.61 0
2009/03/10 5,195.52 5,185.63 5,172.92 5,336.58 5,149.11 0
2009/02/18 4,683.00 4,887.07 4,688.39 4,680.31 4,807.46 0
Table 11: Min-Cost Redispatch Costs: Nov 23, Mar 10, Feb 18
@@
On 2009/11/23 and 2009/03/10, the allocation obtained from the Euphemia algorithm is identical to the national configuration, and no welfare loss occurs. On 2009/02/18, costs under the Euphemia algorithm are slightly higher by 0.36%. This validates our assertion that differences between zonal models – including Euphemia – are minor and nuanced, while nodal pricing constitutes a more substantial change in generation and redispatch costs.
@@
2009/11/23 31,734.14 31,734.14 31,903.40 31,731.23 31,627.79 29,596.10
2009/03/10 36,597.17 36,679.94 36,667.30 36,998.31 36,796.13 35,200.69
2009/02/18 50,970.23 51,358.40 50,975.96 50,967.65 51,240.14 49,303.25
Table 12: Total Costs: Nov 25, Mar 10, Feb 18
Figure 7 illustrates the locational IP, CH, and Join prices for the three selected days, averaged over 24 hours. As previously discussed, the prices on these days exhibit a positive correlation with generation costs, with the exception of Northern Central and Southern Germany with consistent price outliers due to local supply structures. Notably, even though at lower price levels, a more distinct north-south price gradient can be observed on 2009/11/23 compared to 2009/03/10. Specifically, on 2009/11/23, demand is low nationwide, yet inexpensive electricity from the north cannot be transmitted to the south, leading to uneven price levels. In contrast, on 2009/03/10, with higher demand but a balanced grid, prices are uniform across the country. This phenomenon holds for all three pricing rules. The simplified zonal models fail to detect any transmission bottlenecks on 2009/11/23, i.e., cross-zonal constraints are not tight and there are no price variations between zones. Consequently, if the zone configurations are suboptimal, zonal prices do not signal scarcity appropriately.
@@
(d) 2009/03/10 - Avg IP Prices
(e) 2009/03/10 - Avg CH Prices
(f) 2009/03/10 - Avg Join Prices